0Let X1, ....., Xn be iid Random variable from a Uniform distribution (-\theta, \theta) with pdf given byf_x(x|\theta) = \frac{1}{2\theta}*I(-\theta < x < \theta), \theta > 0 .

(1) Is the 2-dimensional statistics T1(X) = (X(1), X(n)) a complete sufficient statistics? Justify your answer

(2) Is the one-dimensional statistic T_2(x) = max_i\{|X_i|\} a complete sufficient statistic? Justify your answer
Answer #1

(1) solution :

Given data :

2-dimensional statistics T1(X) = (X(1)

Now we have to justify :

Is the 2-dimensional statistics T1(X) = (X(1), X(n)) a complete sufficient statistics:

Yes, the 2-dimensional statistics T1(X) = (X(1), X(n)) is a complete sufficient statistics.

  • The given two dimensional measurement is a finished adequate measurement
  • Since we can get the unbiased assessed estimation of the main parameter

(2) solution :

Given data :

one-dimensional statistic  

Now, we have to justify :

Is the one-dimensional statistic a complete sufficient statistic:

Yes,  the one-dimensional statistic a complete sufficient statistic

  • The given one dimensional measurement is a finished adequate measurement
  • Since we can get the unprejudiced assessed estimation of the main parameter .

Note: As per the HOMEWORKLIB RULES, two questions are enough. so i am answered the question 1 and 2 .If you want remaining  question please re upload as another question.

Thank you,